Description: | Below is an overview from Chris Fiebrich, the Mesonet QA manager
of the automated quality assurance results for the March 2001
Oklahoma Mesonet data.
69 trouble tickets were resolved by the techs during March.
WOW! You won't notice all of that work here in this report,
because some of the fixes were for scheduled rotations of
sensors and work done on OASIS sensors (both of which we
haven't made a part of the monthly QA report as of yet.)
Soil temperature probes under bare soil continue to be a
major problem as over 20 stations have been found to have
their probes literally heaved out of the ground. This is
something we've never witnessed in the Mesonet EVER, and
is likely due to all the wet, freezing weather we had this
past winter. We'll hope it doesn't happen again for at least
8 more years.
The Mesonet technicians are beginning their Spring Pass 2001
in which all sites will be visited over the next 3 months.
During this pass, they will be taking site photos, maintaining
the vegetation inside the site enclosure, cleaning sensors,
and performing rain gauge tests.
===============================================================
TAIR Current: None
Added: None
Resolved: #5442 GRA2 Replaced faulty sensor
RELH Current: None
Added: None
Resolved: #5505 HUGO Replaced lightning-damaged
sensor
WDIR Current: None
Added: None
Resolved: None
WSPD Current: None
Added: None
Resolved: #5490 RING Replaced damaged sensor
PRES Current: None
Added: None
Resolved: None
SRAD Current: #5519 MEDF Monthly QA indicates 10% low
bias
Added: None
Resolved: #5072 KETC Replaced sensor with high bias
#5159 FAIR Replaced lightning-damaged
sensor
#5394 SEIL Replaced sensor with low bias
RAIN Current: #5511 HUGO Gauge occasionally double-tips
Added: None
Resolved: #5397 MEDF Replaced bad switch
#5466 PUTN Replaced bad switch
TA9M Current: None
Added: None
Resolved: #5492 JAYX Repaired lose wire
WS2M Current: None
Added: None
Resolved: None
TS10 Current: None
Added: None
Resolved: #5507 HUGO Replaced lightning-damaged
sensor
TB10 Current: #5518 MEDF Monthly QA indicates 3 C warm
bias
Added: None
Resolved: #5402 MAYR Replaced bad sensor
#5468 STUA Replaced gopher-damaged sensor
#5440 WILB Corrected erosion problem
TS05 Current: #5194 FAIR 5 C cool bias
Added: None
Resolved: #5508 HUGO Replaced lightning-damaged sensor
TB05 Current: #5515 STIG Sensor reporting erratic values
Added: #5569 CATO Check for erosion
#5570 BREC Check for erosion
Resolved: #5419 MARS Repaired erosion problem
#5454 BLAC Repaired erosion problem
#5455 KING Repaired erosion problem
#5456 OKMU Repaired erosion problem
#5457 PRYO Repaired erosion problem
#5458 FORA Site checked for erosion
#5459 TAHL Repaired erosion problem
#5499 TISH Repaired erosion problem
#5535 BIXB Repaired erosion problem
#5536 WYNO Repaired erosion problem
#5537 COPA Repaired erosion problem
#5538 NOWA Repaired erosion problem
#5539 MIAM Repaired erosion problem
#5541 WILB Repaired erosion problem
TS30 Current: #5517 EUFA Monthly QA indicates 3 C cool
bias
Added: None
Resolved: #5509 HUGO Replaced lightning-damaged
sensor
----------------------------------------------------------------
ARS QA REPORT
----------------------------------------------------------------
TAIR Current: None
Added: None
Resolved: #5522 A152 Repaired vandalized sensor
RELH Current: None
Added: None
Resolved: #5516 A149 Replaced sensor with high
bias
SRAD Current: None
Added: None
Resolved: None
RAIN Current: None
Added: None
Resolved: #5467 A121 Replaced bad switch
TS05 Current: None
Added: None
Resolved: #5453 A165 Checked for erosion
TS10 Current: None
Added: None
Resolved: None
TS15 Current: None
Added: None
Resolved: #5523 A165 Corrected wiring problem
TS30 Current: None
Added: None
Resolved: #5524 A165 Corrected wiring problem
========================================================================
LEGEND
TAIR Air Temperature measured at 1.5 meters
RELH Relative Humidity measured at 1.5 meters
WDIR Wind Direction measured at 10 meters
WSPD Wind Speed measured at 10 meters
PRES Pressure
SRAD Incident Solar Radiation
RAIN Rainfall
TA9M Air Temperature measured at 9 meters
WS2M Wind Speed measured at 2 meters
TS10 Soil Temperature measured at 10 cm under native sod
TB10 Soil Temperature measured at 10 cm under bare soil
TS05 Soil Temperature measured at 5 cm under native sod
TB05 Soil Temperature measured at 5 cm under bare soil
TS15 Soil Temperature measured at 15 cm under native sod
TS30 Soil Temperature measured at 30 cm under native sod
======================================================================= |